# Testing Stochastic Block Models Based on Maximum Sampling Entry-Wise Deviations

### v1

@@ 2.4 The asymptotic power
This condition is mild and analogous to Condition (C2). Basically, it imposes bounds on community sizes for the hypothetical number of communities $K_{0}$ and the hypothetical membership vector $g_{0}$. Consider the test statistic $T_{n,2}=F_{n}^{2}(K_{0},g_{0})-2\log(MK_{0})+\log\log(MK_{0})$ defined in Subsection 2.3. We next demonstrate its asymptotic property under alternatives.
@@
$P\Big{\{}T_{n,2}\geq c_{1}\log(MK_{0})\Big{\}}\to 1.$
Theorem 2 indicates that $T_{n,2}$ diverges at a rate of $\log(MK_{0})$ under the alternative hypothesis in (II) that $(K_{0},g_{0})\in\mathcal{D}(K,g,Q)$. This result is analogous to Theorem 2 in Hu et al. (2021). The main difference is that our proposed $Q$ and $Q^{g_{0}}$ can decay to zero in sparse networks under the alternative set $\mathcal{D}(K,g,Q)$. Accordingly, the resulting test statistic is applicable for both sparse and dense networks. It is worth noting that Theorem 2 can also be used to calculate the power of $T_{n,1}$ under the alternative hypothesis in (I).
Under the null hypothesis in (II), Theorem 1 shows $T_{n,2}=O_{p}(1)$. In contrast, under alternatives, Theorem 2 indicates that the divergence rate of $T_{n,2}$ is $\log(MK_{0})$. Then, applying Condition (C3)-(iii), $MK=O(n^{\xi_{4}})$ with $10<\xi_{4}<$ 1, the rate of $T_{n,2}$ is $\log(n)$, which is the same as that of Hu et al. (2021). From a theoretical perspective, the test statistic $T_{n,2}$, with divergence rate $\log(n)$, is sufficient to separate the null and alternatives asymptotically. In addition, under Conditions (C3)-(ii), $K=o\{(n/\log n)^{\xi_{3}}\}$ with $20<\xi_{3}<1/$ 2, and (C3)-(iii), $MK=O(n^{\xi_{4}})$ with $10<\xi_{4}<$ 1, we have that $M=O(n^{\xi_{4}})$. As a result, $T_{n,2}$ is asymptotically powerful against the alternative set $\mathcal{D}(K,g,Q)$, even for moderately large $M$.
It is noteworthy that we have provided three sufficient conditions to ensure that the alternative $(K_{0},g_{0})$ is contained in $\mathcal{D}(K,g,Q)$; see Proposition 1 in Section S.4 and its proof in Sections S.5 of the supplementary material. In addition, according to both an AE’s and an anonymous referee’s guidance, we have rigorously established the power function of $T_{n,2}$ under some special alternatives in the set $D(K,g,Q)$, and found that the power of $T_{n,2}$ is expected to increase as the network becomes denser (i.e., $Q$ increases), the network size $n$ grows, or the number of communities $K$ decreases; see Proposition 2 and its related discussion in Section S.4, and the proof in Section S.6 of the supplementary material.
@@ 2.5 Bootstrap-corrected test statistic
Note that the speed at which $T_{n,1}$ or $T_{n,2}$ converges to the extreme distribution may be slow. We next consider a finite-sample bootstrap-corrected test statistic by generating bootstrap samples to improve the finite sample performance. This bootstrap correction was proposed by Bickel and Sarkar (2016) and also considered by Lei (2016) and Hu et al. (2021). In the remainder of this article, we use generic notation $T_{n}$ to denote either $T_{n,1}$ or $T_{n,2}$ for testing (I) or (II). Given an adjacency matrix $A$, $K=K_{0}$ and $g=g_{0}$, the bootstrap-corrected test statistic $T_{n,boot}$ can be obtained using Algorithm 2 below to test (II), which consists of four steps. Note that $g_{0}$ can be replaced by its consistent estimator $\widehat{g}$ for testing (I).
@@
Output: Bootstrap-corrected test statistic $T_{n,boot}$, and the decision in regard to the null hypothesis being rejected or not.
1 Calculate $\widehat{Q}$ and then obtain $T_{n}$ from Algorithm 1 using $A$, $g_{0}$, $K_{0}$, $B$ and $M$.
2For $j=1,\cdots,J$, randomly generate a network $A^{(j)}$ based on the stochastic block model $(\widehat{Q},g_{0})$, and obtain $T^{(j)}_{n}$ from Algorithm 1 using $A^{(j)}$, $(\widehat{Q},g_{0})$, $K_{0}$, $B$ and $M$.
3Estimate the location and scale parameters $\widehat{\mu}$ and $\widehat{\beta}$ of the Type-I extreme distribution using the maximum likelihood estimation method given $T^{(j)}_{n}$s for $j=1,\cdots,J$.
Calculate the bootstrap-corrected test statistic $T_{n,boot}=\mu+\beta(\frac{T_{n}-\widehat{\mu}}{\widehat{\beta}})$, where $\mu=-2\log(\sqrt{\pi})$ and $\beta=2$. Reject the null hypothesis if $T_{n,boot}>q_{\alpha}$.
@@
We can demonstrate theoretically that the bootstrap-corrected test statistic $T_{n,boot}$ has the same asymptotic null distribution as $T_{n}$; see Corollary S.2 and its proof in Section S.7 of the supplementary material. To examine the finite sample performance of these two tests, simulation studies are reported in the next section. They indicate that $T_{n,boot}$ performs better than $T_{n}$ in both dense and sparse networks.
@@ 2.6 An augmented test
As noted by Hu et al. (2021), the power of their proposed test for hypothesis test (I) may be small when the networks are partitioned equally. To improve the power, they proposed a novel technique of adding a new community to the observed network. Consider a planted partition model from an adjacency matrix $A$ with $n$ nodes and $K$ equal-sized communities. Let $K^{+}_{0}=K_{0}+1$. Inspired by the approach of Hu et al. (2021), we propose an augmented test statistic to test (I). In Algorithm 3 below, we first obtain the test statistic $T^{+}_{n}$ to test (II), which consists of seven steps. Then we replace $g_{0}$ in $T^{+}_{n}$ with its consistent estimator $\widehat{g}$ so that it can be used to test (I).
@@
Output: Augmented test statistic $T^{+}_{n}$ and the decision in regard to the null hypothesis being rejected or not.
@@
6At the $m$-th realization for $m=1,\cdots M$, randomly draw $B$ samples from $\{\widehat{\rho}^{+}_{1v},\cdots,\widehat{\rho}^{+}_{n^{+}v}\}$ obtained in Step 4 for each $1\leq v\leq K^{+}_{0}$, and calculate $\widehat{\gamma}^{+}_{mv}$ using (2.3).
7Find the maximum value of $\widehat{\gamma}_{mv}$ for $1\leq m\leq M$ and $1\leq v\leq K^{+}_{0}$ and obtain $F_{n}(K^{+}_{0},g_{0}^{+})$ using (2.4).
Calculate augmented test statistic $T^{+}_{n}=F^{2}_{n}(K^{+}_{0},g_{0}^{+})-2\log(MK^{+}_{0})+2\log\log(MK^{+}_{0})$ and reject the null hypothesis if $T^{+}_{n}>q_{\alpha}$.
@@
It is worth noting that the above augmented procedure in Step 2 yields a larger disparity measurement, denoted by $q_{vu^{*}}$ in equation (2.7). As a result, $\sqrt{B}\max_{1\leq v\leq K_{0}}q_{v}/\sqrt{\log(MK_{0})}$ becomes larger, which leads to a larger value of $T^{+}_{n}$ so that its power is improved. Additionally, we can demonstrate theoretically that the asymptotic null distribution of $T^{+}_{n}$ is identical to that of $T_{n}$; see Corollary S.3 in Section S.7 of the supplementary material. Moreover, to improve the power of the test statistic by controlling the type-I error close to the 5% nominal level, we can integrate the bootstrap-corrected method introduced in Subsection 2.5 with this augmented test statistic and obtain the augmented bootstrap-corrected statistic $T^{+}_{n,boot}$. Analogously, one can verify that the asymptotic null distribution of $T^{+}_{n,boot}$ is the same as that of $T_{n}$. According to the simulation results, $T^{+}_{n,boot}$ can not only control the size well, but also improve the power. Hence, we include $T^{+}_{n,boot}$ in our simulation studies.
@@ 3 Simulation studies
In studying hypothesis test (I), we not only consider our method, but also include the four competing methods proposed by Hu et al. (2021), Han et al. (2023), Zhang and Amini (2023), and Lei (2016). After including the augmented or the bootstrap-corrected procedure applied to our method, Hu et al. (2021)’s method, and Lei (2016)’s method, we include a total of eight tests in our simulation studies for testing (I). They are $T_{n}$, $T^{+}_{n,boot}$, $T_{Hu}$, $T^{+}_{Hu,boot}$, $T_{Lei}$, $T_{Lei,boot}$, $T_{Han}$, and $T_{Zhang}$, where $T_{Hu}$ and $T^{+}_{Hu,boot}$ are the test statistics proposed by Hu et al. (2021). It is worth noting that the augmented procedure is proposed to test (I), and the methods of Lei (2016), Han et al. (2023), and Zhang and Amini (2023) are not applicable to test (II). Accordingly, we only consider the test statistics $T_{n}$ and $T_{Hu}$ in testing (II). Due to space constraint, all the simulation results for hypothesis test (II) are given in the supplementary material; see Tables S.1-S.2 of Section S.8. In simulation studies, all simulation results are evaluated using 500 realizations.
@@ 3.1 Simulations under dense networks
Tables 1-2 present the sizes and powers of all eight tests for cases (i) and (ii), respectively. They show the following findings. First, our proposed tests $T_{n}$ and $T^{+}_{n,boot}$ generally perform better than $T_{Hu}$ and $T^{+}_{Hu,boot}$ in controlling the size, and the sizes of $T_{Hu}$ and $T^{+}_{Hu,boot}$ are significantly and slightly distorted, respectively, when $K>10$. This finding is reasonable, because the sampling process can mitigate the instability of the spectral clustering estimate $\widehat{g}$. Recall that our proposed statistic is constructed by sampling the $B$ entry-wise deviations obtained from Hu et al. (2021). Hence, as $B\rightarrow\infty$, the negative impact due to a small number of poorly estimated $\widehat{g}$ can be ignored. Thus, the performance of the proposed statistic can be dominated by a large number of consistently estimated membership vectors $\widehat{g}$. In addition, both $T_{n}$ and $T^{+}_{n,boot}$ outperform $T_{Lei}$, $T_{Lei,boot}$, $T_{Han}$ and $T_{Zhang}$ in controlling the size. Specifically, the latter four methods exhibit significant size distortions when $K\geq 8$, since their empirical sizes are significantly larger than the nominal level of 5%. Consequently, $T^{+}_{n,boot}$ and $T_{n}$ perform the best overall for controlling the size in dense networks.
@@
Second, both $T_{n}$ and $T_{Hu}$ exhibit weaker power against alternatives than those of $T_{Lei}$, $T_{Lei,boot}$, $T_{Han}$, and $T_{Zhang}$ in dense networks, and $T_{n}$ is also less powerful than $T_{Hu}$. As expected, the sizes and powers of $T_{n}$ and $T_{Hu}$ can be enhanced by $T^{+}_{n,boot}$ and $T^{+}_{Hu,boot}$, respectively, since they integrate the bootstrap-corrected method with the augmentation procedure. Tables 1-2 demonstrate that the empirical powers of $T^{+}_{n,boot}$ and $T^{+}_{Hu,boot}$ are comparable to those of $T_{Lei}$, $T_{Lei,boot}$, and $T_{Zhang}$, and stronger than that of $T_{Han}$. In addition, $T^{+}_{n,boot}$ is less powerful than $T^{+}_{Hu,boot}$, in particular when $K$ is large, as seen in Table 2. This finding is not surprising since the size of $T^{+}_{Hu,boot}$ is distorted for large $K$. Overall, $T^{+}_{n,boot}$ is a powerful test.
@@
In sum, Tables 1-2 demonstrate that $T^{+}_{n,boot}$ performs the best overall as compared with other tests in testing (I), and $T_{n}$ controls size well.
@@ 3.2 Simulations under sparse networks
Based on the network density conditions imposed by Lei (2016), Hu et al. (2021), Zhang and Amini (2023), and Han et al. (2023), their resulting test statistics, $T_{Lei}$, $T_{Lei,boot}$, $T_{Hu}$, $T^{+}_{Hu,boot}$, $T_{Zhang}$, and $T_{Han}$, are not suitable for networks with sparse density on the order of $O(n^{-1}\log\log n)$. In addition, the methods proposed by Zhang and Amini (2023) and Han et al. (2023) required a fixed value of $K$. Hence, their test statistics may not be applicable to data with a large $K$. Our simulation results in Tables 3–4 corroborate these assertions. Specifically, the empirical sizes of $T_{Lei}$, $T_{Hu}$, and $T^{+}_{Hu,boot}$ are all significantly greater than 5% and suffer from size distortions. Additionally, the empirical sizes of $T_{Zhang}$ are somewhat erratic and none of them approximate to the 5% significance level. As for the tests $T_{Han}$ and $T_{Lei,boot}$, they perform poorly for $K\geq 3$, exhibiting size distortions. On the other hand, the sizes of $T_{n}$ and $T^{+}_{n,boot}$ are close to the nominal level of 5%. Since the sizes of the six competing tests are distorted, it is not sensible to make power comparisons. In sum, under sparse networks, Tables 3-4 indicate that our proposed tests, $T_{n}$ and $T^{+}_{n,boot}$, perform well in comparison with the other six tests, and $T^{+}_{n,boot}$ is more powerful than $T_{n}$, which supports our theoretical results.
Based on Tables 1-4, we conclude that our proposed test $T^{+}_{n,boot}$ performs the best overall in testing (I) under both dense and sparse networks, and $T_{n}$ can control empirical sizes well. In addition, Hu et al. (2021)’s test, $T^{+}_{Hu,boot}$, can be used for dense networks when $K$ is not large.
@@
Remark 3. We also construct simulation studies under Settings 1 and 2 to test hypothesis (II). Due to space constraint, all of the simulation results for testing hypothesis (II) are given in Tables S.1-S.2 of Section S.8 in the supplementary material. The simulation results in Tables S.1-S.2 are consistent with those findings in Tables 1-4. Under Setting 1, Table S.1 indicates that our proposed statistic $T_{n}$ is superior to $T_{Hu}$ in the size control, and both $T_{n}$ and $T_{Hu}$ are powerful against the alternatives with the empirical powers all equal to 1. Under Setting 2, Table S.2 shows that the empirical sizes of $T_{n}$ are close to the nominal level of 5% and the empirical powers of $T_{n}$ are all 1, while $T_{Hu}$ fails to control size well. In sum, when testing hypothesis (II), $T_{n}$ performs well and is superior to $T_{Hu}$ in the size control under both dense and sparse networks.
Remark 4. It is worth noting that the condition $D_{n}=O(n^{-1}\log n)$ has been considered in obtaining theoretical properties of various tests; see, e.g., the $T_{Zhang}$ test. According to this condition, we construct additional simulation studies under Setting 2 to test hypotheses (I) and (II), except that the edge probabilities between any two communities $u$ and $v$ are adjusted to $2\log n(1+4\times I(u=v))/n$ and $3\log n(1+4\times I(u=v))/n$, respectively, for cases (i) and (ii). Additional simulation results under this setting are presented in Tables S.3-S.5 of Section S.9 in the supplementary material. The simulation results in Tables S.3-S.5 are consistent with those findings in Tables 3-4. In sum, when testing hypothesis (I), our proposed test statistics $T_{n}$ and $T^{+}_{n,boot}$ outperform all competing methods, and $T^{+}_{n,boot}$ is more powerful than $T_{n}$ and performs the best overall. When testing hypothesis (II), $T_{n}$ is superior to $T_{Hu}$.

### v2

@@ 2.4 The asymptotic power
This condition is mild and analogous to Condition (C2). Basically, it imposes bounds on community sizes for the hypothetical number of communities $K_{0}$ and the hypothetical membership vector $g_{0}$. Consider the test statistic $\Theta_{n,2}=\Gamma_{n}^{2}(K_{0},g_{0})-2\log(MK_{0})+\log\log(MK_{0})$ defined in Subsection 2.3. We next demonstrate its asymptotic property under alternatives.
@@
$P\Big{\{}\Theta_{n,2}\geq c_{1}\log(MK_{0})\Big{\}}\to 1.$
Theorem 2 indicates that $\Theta_{n,2}$ diverges at a rate of $\log(MK_{0})$ under the alternative hypothesis in (II) that $(K_{0},g_{0})\in\mathcal{D}(K,g,Q)$. This result is analogous to Theorem 2 in Hu et al. (2021). The main difference is that our proposed $Q$ and $Q^{g_{0}}$ can decay to zero in sparse networks under the alternative set $\mathcal{D}(K,g,Q)$. Accordingly, the resulting test statistic is applicable for both sparse and dense networks. It is worth noting that Theorem 2 can also be used to calculate the power of $\Theta_{n,1}$ under the alternative hypothesis in (I).
Under the null hypothesis in (II), Theorem 1 shows $\Theta_{n,2}=O_{p}(1)$. In contrast, under alternatives, Theorem 2 indicates that the divergence rate of $\Theta_{n,2}$ is $\log(MK_{0})$. Then, applying Condition (C3)-(iii), $MK=O(n^{\xi_{4}})$ with $10<\xi_{4}<$ 1, the rate of $\Theta_{n,2}$ is $\log(n)$, which is the same as that of Hu et al. (2021). From a theoretical perspective, the test statistic $\Theta_{n,2}$, with divergence rate $\log(n)$, is sufficient to separate the null and alternatives asymptotically. In addition, under Conditions (C3)-(ii), $K=o\{(n/\log n)^{\xi_{3}}\}$ with $20<\xi_{3}<1/$ 2, and (C3)-(iii), $MK=O(n^{\xi_{4}})$ with $10<\xi_{4}<$ 1, we have that $M=O(n^{\xi_{4}})$. As a result, $\Theta_{n,2}$ is asymptotically powerful against the alternative set $\mathcal{D}(K,g,Q)$, even for moderately large $M$.
It is noteworthy that we have provided three sufficient conditions to ensure that the alternative $(K_{0},g_{0})$ is contained in $\mathcal{D}(K,g,Q)$; see Proposition 1 in Section S.4 and its proof in Sections S.5 of the supplementary material. In addition, we have rigorously established the power function of $\Theta_{n,2}$ under some special alternatives in the set $D(K,g,Q)$, and found that the power of $\Theta_{n,2}$ is expected to increase as the network becomes denser (i.e., $Q$ increases), the network size $n$ grows, or the number of communities $K$ decreases; see Proposition 2 and its related discussion in Section S.4, and the proof in Section S.6 of the supplementary material.
@@ 2.5 Bootstrap-corrected test statistic
Note that the speed at which $\Theta_{n,1}$ or $\Theta_{n,2}$ converges to the extreme distribution may be slow. We next consider a finite-sample bootstrap-corrected test statistic by generating bootstrap samples to improve the finite sample performance. This bootstrap correction was proposed by Bickel and Sarkar (2016) and also considered by Lei (2016) and Hu et al. (2021). In the remainder of this article, we use generic notation $\Theta_{n}$ to denote either $\Theta_{n,1}$ or $\Theta_{n,2}$ for testing (I) or (II). Given an adjacency matrix $A$, $K=K_{0}$ and $g=g_{0}$, the bootstrap-corrected test statistic $\Theta_{n,boot}$ can be obtained using Algorithm 2 below to test (II), which consists of four steps. Note that $g_{0}$ can be replaced by its consistent estimator $\widehat{g}$ for testing (I).
@@
Output: Bootstrap-corrected test statistic $\Theta_{n,boot}$, and the decision in regard to the null hypothesis being rejected or not.
1 Calculate $\widehat{Q}$ and then obtain $\Theta_{n}$ from Algorithm 1 using $A$, $g_{0}$, $K_{0}$, $B$ and $M$.
2For $j=1,\cdots,J$, randomly generate a network $A^{(j)}$ based on the stochastic block model $(\widehat{Q},g_{0})$, and obtain $\Theta^{(j)}_{n}$ from Algorithm 1 using $A^{(j)}$, $(\widehat{Q},g_{0})$, $K_{0}$, $B$ and $M$.
3Estimate the location and scale parameters $\widehat{\mu}$ and $\widehat{\beta}$ of the Type-I extreme distribution using the maximum likelihood estimation method given $\Theta^{(j)}_{n}$s for $j=1,\cdots,J$.
Calculate the bootstrap-corrected test statistic $\Theta_{n,boot}=\mu+\beta(\frac{\Theta_{n}-\widehat{\mu}}{\widehat{\beta}})$, where $\mu=-2\log(\sqrt{\pi})$ and $\beta=2$. Reject the null hypothesis if $\Theta_{n,boot}>q_{\alpha}$.
@@
We can demonstrate theoretically that the bootstrap-corrected test statistic $\Theta_{n,boot}$ has the same asymptotic null distribution as $\Theta_{n}$; see Corollary S.2 and its proof in Section S.7 of the supplementary material. To examine the finite sample performance of these two tests, simulation studies are reported in the next section. They indicate that $\Theta_{n,boot}$ performs better than $\Theta_{n}$ in both dense and sparse networks.
@@ 2.6 An augmented test
As noted by Hu et al. (2021), the power of their proposed test for hypothesis test (I) may be small when the networks are partitioned equally. To improve the power, they proposed a novel technique of adding a new community to the observed network. Consider a planted partition model from an adjacency matrix $A$ with $n$ nodes and $K$ equal-sized communities. Let $K^{+}_{0}=K_{0}+1$. Inspired by the approach of Hu et al. (2021), we propose an augmented test statistic to test (I). In Algorithm 3 below, we first obtain the test statistic $\Theta^{+}_{n}$ to test (II), which consists of seven steps. Then we replace $g_{0}$ in $\Theta^{+}_{n}$ with its consistent estimator $\widehat{g}$ so that it can be used to test (I).
@@
Output: Augmented test statistic $\Theta^{+}_{n}$ and the decision in regard to the null hypothesis being rejected or not.
@@
6At the $m$-th realization for $m=1,\cdots M$, randomly draw $B$ samples from $\{\widehat{\rho}^{+}_{1v},\cdots,\widehat{\rho}^{+}_{n^{+}v}\}$ obtained in Step 4 for each $1\leq v\leq K^{+}_{0}$, and calculate $\widehat{\psi}^{+}_{mv}$ using (2.3).
7Find the maximum value of $\widehat{\psi}^{+}_{mv}$ for $1\leq m\leq M$ and $1\leq v\leq K^{+}_{0}$ and obtain $\Gamma_{n}(K^{+}_{0},g_{0}^{+})$ using (2.4).
Calculate augmented test statistic $\Theta^{+}_{n}=\Gamma^{2}_{n}(K^{+}_{0},g_{0}^{+})-2\log(MK^{+}_{0})+2\log\log(MK^{+}_{0})$ and reject the null hypothesis if $\Theta^{+}_{n}>q_{\alpha}$.
@@
It is worth noting that the above augmented procedure in Step 2 yields a larger disparity measurement, denoted by $q_{vu^{*}}$ in equation (2.7). As a result, $\sqrt{B}\max_{1\leq v\leq K_{0}}q_{v}/\sqrt{\log(MK_{0})}$ becomes larger, which leads to a larger value of $\Theta^{+}_{n}$ so that its power is improved. Additionally, we can demonstrate theoretically that the asymptotic null distribution of $\Theta^{+}_{n}$ is identical to that of $\Theta_{n}$; see Corollary S.3 in Section S.7 of the supplementary material. Moreover, to improve the power of the test statistic by controlling the type-I error close to the 5% nominal level, we can integrate the bootstrap-corrected method introduced in Subsection 2.5 with this augmented test statistic and obtain the augmented bootstrap-corrected statistic $\Theta^{+}_{n,boot}$. Analogously, one can verify that the asymptotic null distribution of $\Theta^{+}_{n,boot}$ is the same as that of $\Theta_{n}$. According to the simulation results, $\Theta^{+}_{n,boot}$ can not only control the size well, but also improve the power. Hence, we include $\Theta^{+}_{n,boot}$ in our simulation studies.
@@ 3 Simulation studies
In studying hypothesis test (I), we not only consider our method, but also include the four competing methods proposed by Hu et al. (2021), Han et al. (2023), Zhang and Amini (2023), and Lei (2016). After including the augmented or the bootstrap-corrected procedure applied to our method, Hu et al. (2021)’s method, and Lei (2016)’s method, we include a total of eight tests in our simulation studies for testing (I). They are $\Theta_{n}$, $\Theta^{+}_{n,boot}$, $\Theta_{Hu}$, $\Theta^{+}_{Hu,boot}$, $\Theta_{Lei}$, $\Theta_{Lei,boot}$, $\Theta_{Han}$, and $\Theta_{Zhang}$, where $\Theta_{Hu}$ and $\Theta^{+}_{Hu,boot}$ are the test statistics proposed by Hu et al. (2021). It is worth noting that the augmented procedure is proposed to test (I), and the methods of Lei (2016), Han et al. (2023), and Zhang and Amini (2023) are not applicable to test (II). Accordingly, we only consider the test statistics $\Theta_{n}$ and $\Theta_{Hu}$ in testing (II). Due to space constraint, all the simulation results for hypothesis test (II) are given in the supplementary material; see Tables S.1-S.2 of Section S.8. In simulation studies, all simulation results are evaluated using 500 realizations.
@@ 3.1 Simulations under dense networks
Tables 1-2 present the sizes and powers of all eight tests for cases (i) and (ii), respectively. They show the following findings. First, our proposed tests $\Theta_{n}$ and $\Theta^{+}_{n,boot}$ generally perform better than $\Theta_{Hu}$ and $\Theta^{+}_{Hu,boot}$ in controlling the size, and the sizes of $\Theta_{Hu}$ and $\Theta^{+}_{Hu,boot}$ are significantly and slightly distorted, respectively, when $K>10$. This finding is reasonable, because the sampling process can mitigate the instability of the spectral clustering estimate $\widehat{g}$. Recall that our proposed statistic is constructed by sampling the $B$ entry-wise deviations obtained from Hu et al. (2021). Hence, as $B\rightarrow\infty$, the negative impact due to a small number of poorly estimated $\widehat{g}$ can be ignored. Thus, the performance of the proposed statistic can be dominated by a large number of consistently estimated membership vectors $\widehat{g}$. In addition, both $\Theta_{n}$ and $\Theta^{+}_{n,boot}$ outperform $\Theta_{Lei}$, $\Theta_{Lei,boot}$, $\Theta_{Han}$ and $\Theta_{Zhang}$ in controlling the size. Specifically, the latter four methods exhibit significant size distortions when $K\geq 8$, since their empirical sizes are significantly larger than the nominal level of 5%. Consequently, $\Theta^{+}_{n,boot}$ and $\Theta_{n}$ perform the best overall for controlling the size in dense networks.
Second, both $\Theta_{n}$ and $\Theta_{Hu}$ exhibit weaker power against alternatives than those of $\Theta_{Lei}$, $\Theta_{Lei,boot}$, $\Theta_{Han}$, and $\Theta_{Zhang}$ in dense networks, and $\Theta_{n}$ is also less powerful than $\Theta_{Hu}$. As expected, the sizes and powers of $\Theta_{n}$ and $\Theta_{Hu}$ can be enhanced by $\Theta^{+}_{n,boot}$ and $\Theta^{+}_{Hu,boot}$, respectively, since they integrate the bootstrap-corrected method with the augmentation procedure. Tables 1-2 demonstrate that the empirical powers of $\Theta^{+}_{n,boot}$ and $\Theta^{+}_{Hu,boot}$ are comparable to those of $\Theta_{Lei}$, $\Theta_{Lei,boot}$, and $\Theta_{Zhang}$, and stronger than that of $\Theta_{Han}$. In addition, $\Theta^{+}_{n,boot}$ is less powerful than $\Theta^{+}_{Hu,boot}$, in particular when $K$ is large, as seen in Table 2. This finding is not surprising since the size of $\Theta^{+}_{Hu,boot}$ is distorted for large $K$. Overall, $\Theta^{+}_{n,boot}$ is a powerful test.
@@
In sum, Tables 1-2 demonstrate that $\Theta^{+}_{n,boot}$ performs the best overall as compared with other tests in testing (I), and $\Theta_{n}$ controls size well.
@@ 3.2 Simulations under sparse networks
Based on the network density conditions imposed by Lei (2016), Hu et al. (2021), Zhang and Amini (2023), and Han et al. (2023), their resulting test statistics, $\Theta_{Lei}$, $\Theta_{Lei,boot}$, $\Theta_{Hu}$, $\Theta^{+}_{Hu,boot}$, $\Theta_{Zhang}$, and $\Theta_{Han}$, are not suitable for networks with sparse density on the order of $O(n^{-1}\log\log n)$. In addition, the methods proposed by Zhang and Amini (2023) and Han et al. (2023) required a fixed value of $K$. Hence, their test statistics may not be applicable to data with a large $K$. Our simulation results in Tables 3–4 corroborate these assertions. Specifically, the empirical sizes of $\Theta_{Lei}$, $\Theta_{Hu}$, and $\Theta^{+}_{Hu,boot}$ are all significantly greater than 5% and suffer from size distortions. Additionally, the empirical sizes of $\Theta_{Zhang}$ are somewhat erratic and none of them approximate to the 5% significance level. As for the tests $\Theta_{Han}$ and $\Theta_{Lei,boot}$, they perform poorly for $K\geq 3$, exhibiting size distortions. On the other hand, the sizes of $\Theta_{n}$ and $\Theta^{+}_{n,boot}$ are close to the nominal level of 5%. Since the sizes of the six competing tests are distorted, it is not sensible to make power comparisons. In sum, under sparse networks, Tables 3-4 indicate that our proposed tests, $\Theta_{n}$ and $\Theta^{+}_{n,boot}$, perform well in comparison with the other six tests, and $\Theta^{+}_{n,boot}$ is more powerful than $\Theta_{n}$, which supports our theoretical results.
Based on Tables 1-4, we conclude that our proposed test $\Theta^{+}_{n,boot}$ performs the best overall in testing (I) under both dense and sparse networks, and $\Theta_{n}$ can control empirical sizes well. In addition, Hu et al. (2021)’s test, $\Theta^{+}_{Hu,boot}$, can be used for dense networks when $K$ is not large.
@@
Remark 4.
We also construct simulation studies under Settings 1 and 2 to test hypothesis (II). Due to space constraint, all of the simulation results for testing hypothesis (II) are given in Tables S.1-S.2 of Section S.8 in the supplementary material. The simulation results in Tables S.1-S.2 are consistent with those findings in Tables 1-4. Under Setting 1, Table S.1 indicates that our proposed statistic $\Theta_{n}$ is superior to $\Theta_{Hu}$ in the size control, and both $\Theta_{n}$ and $\Theta_{Hu}$ are powerful against the alternatives with the empirical powers all equal to 1. Under Setting 2, Table S.2 shows that the empirical sizes of $\Theta_{n}$ are close to the nominal level of 5% and the empirical powers of $\Theta_{n}$ are all 1, while $\Theta_{Hu}$ fails to control size well. In sum, when testing hypothesis (II), $\Theta_{n}$ performs well and is superior to $\Theta_{Hu}$ in the size control under both dense and sparse networks.
Remark 5.
It is worth noting that the condition $D_{n}=O(n^{-1}\log n)$ has been considered in obtaining theoretical properties of various tests; see, e.g., the $\Theta_{Zhang}$ test. According to this condition, we construct additional simulation studies under Setting 2 to test hypotheses (I) and (II), except that the edge probabilities between any two communities $u$ and $v$ are adjusted to $2\log n(1+4\times I(u=v))/n$ and $3\log n(1+4\times I(u=v))/n$, respectively, for cases (i) and (ii). Additional simulation results under this setting are presented in Tables S.3-S.5 of Section S.9 in the supplementary material. The simulation results in Tables S.3-S.5 are consistent with those findings in Tables 3-4. In sum, when testing hypothesis (I), our proposed test statistics $\Theta_{n}$ and $\Theta^{+}_{n,boot}$ outperform all competing methods, and $\Theta^{+}_{n,boot}$ is more powerful than $\Theta_{n}$ and performs the best overall. When testing hypothesis (II), $\Theta_{n}$ is superior to $\Theta_{Hu}$.